# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ E.9 Pre-trained Language Model
We investigate how strongly the ICL performance of RA-DT is influenced by the pre-trained LM used in our domain-agnostic embedding model. In Figure 36, we compare our default choice BERT [Devlin et al., 2019] against four alternative encoder and decoder backbones, namely RoBERTa [Liu et al., 2019], DistilRoBERTa, DistilBERT [Sanh et al., 2019] and DistilGPT2. We find that RA-DT maintains decent performance across all pre-trained LMs, indicating robust retrieval performance across different LMs. Generally, the non-distilled variants outperform their distilled counterparts. Moreover, this experiment suggests a clear advantage of encoder-only models over the decoder-only LM, DistilGPT2. This suggests that the encoder-only LMs are better able to capture the relations between tokens within the token sequence, which leads to more precise retrieval of sub-trajectories and higher down-stream performance.
@@
Figure 36: Effect of the Pre-trained LM. Average performances on Dark-Room 10$\times$10 over the course of training for (a) train and (b) test tasks.
@@ E.10 Effect of $K$ on Algorithm Distillation
Finally, we investigate the effect of $K$ on the performance of AD. $K$ determines the number of episodes that have passed between the current and the context trajectory, which are provided to AD as the context. Consequently, $K$ specifies the extent of improvement observed between subsequent episodes. By default, we use $K=100$ for our experiments on Dark-Room $10\times 10$. Therefore, we conduct an ablation study, in which we very $K$ (see Figure 37. We find that too small values for $K$ (e.g., 1 and 10) result in slow ICL behavior. In contrast, too high values for $K$ (e.g., 500) lead to fast initial progress but suboptimal performance in the long term. Only $K=100$ leads to steady improvement across all interaction episodes. Consequently, AD requires careful tuning of $K$.
Figure 37: Ablation on the number of episodes $K$ in AD that have passed between “current” trajectory and “context” trajectory on Dark-Room 10$\times$10. $K$ determines how much improvement is observed between episodes. We find that performance increases as $K$ increases, but only up to a certain point ($K=100$). With $K=500$, AD improves rapidly in the first few episodes, but then flattens out.
@@ E.11 Convergence of Baselines
In the main experiments on grid-worlds reported in Section 4.1, we found that the baselines AD and DPT only reach sub-optimal performance within the 40 ICL trials. Therefore, we analyse their performance if evaluate for more ICL trials. In Figure 38, we compare the evaluation performance of AD and DPT across 200 ICL trials on the 20 hold-out tasks for Dark-Room $10\times 10$. We find that both method continue to improve towards optimal performance in this environment when given more ICL trials. For this ablation, we found it useful to set $K=50$ in AD (see Appendix E.10) instead of $K=100$ as used in our main experiments over 40 ICL trials.
Figure 38: Evaluation of AD and DPT on Dark-Room 10$\times$10 over 200 ICL trials. Both methods continue to improve towards optimal performance on this environment.

### v3

@@ E.9 Pre-trained Language Model
We investigate how strongly the ICL performance of RA-DT is influenced by the pre-trained LM used in our domain-agnostic embedding model. In Figure 35, we compare our default choice BERT (Devlin et al., 2019) against four alternative encoder and decoder backbones, namely RoBERTa (Liu et al., 2019), DistilRoBERTa, DistilBERT (Sanh et al., 2019), and DistilGPT2. We find that RA-DT maintains decent performance across all pre-trained LMs, indicating robust retrieval performance across different LMs. Generally, the non-distilled variants outperform their distilled counterparts. Moreover, this experiment suggests a clear advantage of encoder-only models over the decoder-only LM, DistilGPT2. This suggests that the encoder-only LMs are better able to capture the relations between tokens within the token sequence, which leads to more precise retrieval of sub-trajectories and higher downstream performance.
@@
Figure 35: Effect of the Pre-trained LM. Average performances on Dark-Room $\times10$ over the course of training for (a) train and (b) test tasks.
@@ E.10 Effect of $K$ on Algorithm Distillation
Finally, we investigate the effect of $K$ on the performance of AD. $K$ determines the number of episodes that have passed between the current and the context trajectory, which are provided to AD as the context. Consequently, $K$ specifies the extent of improvement observed between subsequent episodes. By default, we use $K=100$ for our experiments on Dark-Room $10\times 10$. Therefore, we conduct an ablation study, in which we vary $K$ (see Figure 36. We find that too small values for $K$ (e.g., 1 and 10) result in slow ICL behavior. In contrast, too high values for $K$ (e.g., 500) lead to fast initial progress but suboptimal performance in the long term. Only $K=100$ leads to steady improvement across all interaction episodes. Consequently, AD requires careful tuning of $K$.
Figure 36: Ablation on the number of episodes $K$ in AD that have passed between “current” trajectory and “context” trajectory on Dark-Room $\times10$. $K$ determines how much improvement is observed between episodes. We find that performance increases as $K$ increases, but only up to a certain point ($K=100$). With $K=500$, AD improves rapidly in the first few episodes, but then flattens out.
@@ E.11 Convergence of Baselines
In the main experiments on grid-worlds reported in Section 4.1, we found that the baselines AD and DPT only reach sub-optimal performance within the 40 ICL trials. Therefore, we analyse their performance when evaluating for more ICL trials. In Figure 37, we compare the evaluation performance of AD and DPT across 200 ICL trials on the 20 hold-out tasks for Dark-Room $10\times 10$. We find that both methods continue to improve towards optimal performance in this environment when given more ICL trials. For this ablation, we found it useful to set $K=50$ in AD (see Appendix E.10) instead of $K=100$ as used in our main experiments over 40 ICL trials.
Figure 37: Evaluation of AD and DPT on Dark-Room $\times10$ over 200 ICL trials. Both methods continue to improve towards optimal performance in this environment.